If Pseca + qtana =1 and p2sec2a - q2tan2a =5 then find the value of (9p-2 - 4q-2).

Grade:11

1 Answers

Susmita
5 years ago
pseca+qtana=1
Or,pseca=1-qtana
Take whole square on both side.
p2sec2a-q2tan2a=1-2qtana
or,5=1-2qtana
Or,-2qtana=4
Or,qtana=-2
Or,q=-2/tana
Put qtana=-2 in 1st equation.
pseca=1-qtana=3
Or,p=3/seca
So 9/p2-4/q2
=sec2a-tan2a=1
If it helps please approve.thanks.